# NeuroPilot: A Realtime Brain-Computer Interface system to enhance concentration of students in online learning

### v1

@@ VI-B Graphical User Interface Development
The software’s graphical user interface (GUI) was developed using PyQt6. It displays the raw EEG signal in real time, along with the attention and meditation scores from the Crimson SDK, the proposed algorithm’s prediction, the headband’s battery level, and its contact state. An input box is provided to enter the model of the specific headband, along with a Connect button to establish the connection. After connecting the headband and starting the algorithm, if non-attention is detected for five segments consecutively, an alert sound is played and an on-screen warning reminds the subject to concentrate. If contact with the headband is lost, a similar warning prompts the subject to adjust it. For segments with significant artifacts, a message in the GUI indicates that the signal is noisy. When the Stop button is pressed at the end of a session, a summary is displayed showing detailed statistics of attention and non-attention over time. This ensures interactive communication with the subject while minimizing interventions. Therefore better cognitive engagement with the video is promoted and attention span is improved. The developed graphical user interface has been illustrated in Fig. 5a. Additionally, the working pipeline of the real-time feedback incorporated with online learning environment can be visualized from Fig. 5b. It can be seen that the student is alerted when attention is lost during online learning environments. The overall methodology helps in this feedback and the student regains concentration again.
@@ VI-C Pilot Evaluation
To preliminarily assess the effectiveness of the implemented feedback mechanism, a pilot study was conducted involving five participants. Each participant underwent a 10-minute session where a video stimuli of an educational lecture was given. The session consisted of two consecutive phases: a 5-minute baseline phase without feedback and a 5-minute feedback phase. During the baseline phase, the system continuously monitored cognitive states of the participant without any visual or audio alerts on non-attentive state detection. For the feedback phase, the same monitoring was done and the system issued visual and audio alerts whenever a participant remained non-attentive as per the model for approximately 8.4 consecutive seconds (4 overlapped segments).
A paired t-test was conducted on the mean durations of non-attentive states in the two phases. At first, the durations of continuous non-attentive states have been calculated and segments having less than 8 seconds duration has been discarded as the effect of feedback mechanism is being evaluated. The paired t-test yielded a t-statistic of 5.73, and a p-value of 0.007. Moreover, the mean duration of non-attention periods for the participants has been found to be 37.50 seconds in the baseline phases and 14.97 seconds in the feedback phases. Both analyses indicated a significant improvement in concentration during the feedback phase, suggesting that the real-time alerts effectively supported participants in regaining and sustaining attention. Fig. 5c gives a graphical overview for better visualization, portraying the timeline of sustained non-attention in the baseline phase and the feedback phase.
@@ VII Conclusion
As traditional methods like questionnaire and oberservations require manual interventions, they are not suitable for online learning. For this reason, an automated real-time attention classification with neurofeedback mechanism has been developed in this article. The methodology includes data collection through the FocusCalm headband, data preprocessing including segmentation with a sliding window, filtering, and high amplitude artifact removal, feature extraction and selection, and classification. The proposed algorithm achieved an average accuracy of 88.77% on leave-one-subject-out validation, which proves the robustness of the model when faced with unseen test data.
Future work will focus on expanding the dataset to include a more diverse participant pool. Better and more robust methods will be developed to ensure better performance of the model without requiring the attention and meditation scores of the Crimson SDK. The application will be analyzed through implementation in multiple use cases, followed by systematic result collection. The use of multi-channel, consumer grade and ready-to-use EEG devices will be adopted for better spatial coverage to capture and classify visual, auditory, and task distractions alongside attention and drowsy states in learning environments. Another possible implication is integrating this research with multi-modal inputs such as facial expressions and behavioral logs, to evaluate the key dimensions of a class: learner engagement and content quality. This research thus paves the way for practical and scalable real-time BCI applications in education and beyond.

### v2

@@ VI-B Graphical User Interface Development
The graphical user interface (GUI), developed with PyQt6, displays real-time EEG signals, SDK metrics, the algorithm’s prediction, and headband status. The GUI is shown in Fig. 5a. The system provides real-time feedback: an audio-visual alert is triggered after five consecutive non-attention segments. Similar warnings prompt the user to adjust the headband if contact is lost or if significant artifacts are detected. At the end of a session, a summary screen displays detailed attention statistics. This feedback loop is visualized in Fig. 5b, which aims to promote cognitive engagement.
@@ VI-C Pilot Evaluation
To preliminarily assess the effectiveness of the implemented feedback mechanism, a pilot study was conducted involving five participants. Each participant watched a 10-minute educational video. The video was split in two consecutive phases: a 5-minute baseline phase without feedback and a 5-minute feedback phase. During the baseline phase, the system continuously monitored cognitive states without alerts. For the feedback phase, the same monitoring was done and the system issued visual and audio alerts whenever a participant remained non-attentive as per the model for approximately 8.4 consecutive seconds (4 overlapped segments).
A paired t-test was conducted on the mean durations of non-attentive states in the two phases. At first, non-attentive segments with less than 8 seconds duration were discarded as the effect of feedback mechanism was being evaluated. The paired t-test yielded a t-statistic of 5.73, and a p-value of 0.007. Moreover, the mean duration of non-attention periods for the participants was found to be 37.50 seconds in the baseline phases and 14.97 seconds in the feedback phases. Both analyses indicated a significant improvement in concentration during the feedback phase, suggesting that the real-time alerts effectively supported participants in regaining and sustaining attention. Fig. 5c gives a graphical overview for better visualization, portraying the timeline of sustained non-attention in the baseline phase and the feedback phase.
@@ VII Conclusion
As traditional methods like questionnaire and observations require manual interventions, they are not suitable for online learning. For this reason, an automated real-time attention classification with neurofeedback mechanism has been developed in this article. The methodology, based on data from the FocusCalm headband, employed advanced preprocessing, WPT-based feature extraction, and an SVM classifier. The proposed algorithm achieved an average accuracy of 88.77% on leave-one-subject-out validation. This result demonstrates the robustness of the model on unseen subjects and proves the feasibility of using consumer-grade BCI for real-time attention monitoring.
Despite these promising results, some limitations are acknowledged. The model’s performance was validated on a limited participant pool. Furthermore, the optimal feature set includes proprietary scores from the Crimson SDK, which may limit the model’s transparency and adaptability.
Future work will directly address these limitations. We will first focus on expanding the dataset to include a more diverse participant pool. Better and more robust methods will be developed to ensure better performance of the model without requiring proprietary Crimson SDK scores. The application will be analyzed through implementation in multiple use cases, followed by systematic result collection. The use of multi-channel, consumer-grade and ready-to-use EEG devices will be adopted for better spatial coverage to capture and classify visual, auditory, and task distractions alongside the current attention and non-attention states. Another possible implication is integrating this research with multi-modal inputs such as facial expressions and behavioral logs, to evaluate the key dimensions of a class: learner engagement and content quality. This research thus paves the way for practical and scalable real-time BCI applications.